40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 13 03:10:35 2015, Jenne, Update, LSC, 6+ CARM->REFL transitions, 1 DARM->AS transition AS55whitening_lockloss_12March2015.pdfMoreDARMB_powerWentDown_12March2015.pngcarm_cm_up_zip.sh.gz
    Reply  Fri Mar 13 14:11:59 2015, rana, Update, LSC, 6+ CARM->REFL transitions, 1 DARM->AS transition 
Message ID: 11139     Entry time: Fri Mar 13 03:10:35 2015     Reply to this: 11140
Author: Jenne 
Type: Update 
Category: LSC 
Subject: 6+ CARM->REFL transitions, 1 DARM->AS transition 

Much more success tonight.  I only started my tally after I got the CARM transition to work entirely by script, and I have 6 tally marks, so I probably made the CARM to RF-only transition 7 or maybe 8 times tonight in total.  Unfortunately, I only successfully made the DARM transition to AS55 once.    From the wall striptool, counting the number of times the transmitted power went high, I had about 40 lock trials total. 

The one RF-only lock ended around 1:27am.

I think 2 things were most important in their contributions to tonight's success.  I modified the bounceRoll filters in the CARM and DARM filter banks to eat less phase.  Also, using Q's recipe as inspiration, I started engaging the AO path partway through the CARM transition which makes it much less delicate. 


Bounce roll filter

Koji and I added a ~29Hz resonant gain in the bounce roll filter several months ago, to squish some noise that we were seeing in the CARM and DARM ALS error signals.  This does a lot of the phase-eating.  I'm assuming / hoping that that peak won't be present in the CARM and DARM RF error signals.  But even if it is, we can deal with it later.  For now, that peak is not causing so much motion that I require it.  So, it's gone. 

This allowed me to move the complex zero pair from 30 Hz down to 26 Hz.  Overall I think this gained me about 10 degrees of phase at 100Hz, and moved the low end of the phase bubble down by about 10Hz. 


Prep for REFL 11 I through the CM board and CM_SLOW

In order to use Q's recipe (elog 11138), I wanted to be able to lock CARM on REFL11 using the CM_SLOW filter bank. 

I did a few sweeps through CARM resonance while holding on ALS, and determined that the REFL1 input to the CM board needed a gain of -20dB in order to match the slope of CM_SLOW_OUT to CARM_IN (ALS), leaving all of Q's other settings alone.  Q had been using a REFL1 gain of 0dB for the PRY earlier today.

I needed to flip the sign in the input matrix relative to what Q had (he was using +1 in the CM_SLOW -> CARM_B, I used -1 there).  To match this in the fast path, I flipped the polarity of the CM board (Q was using minus polarity, I am using positive).

The CM_SLOW filter bank had a gain of 0.000189733.  I assume that Q did this so that the input matrix element could be unity.  I left this number alone.  It is of the same order as the plain REFL11I->CARM input matrix element of 1e-4 from Saturday night, so it seemed fine.

During my sweeps through CARM resonance, I also saw that I needed an offset to make CM_SLOW's average about 0.  With the crazy gain number, I needed an offest of -475 in the CM_SLOW filter bank.  As I type this though, it occurs to me that I should have put this in the CM board, since the fast path will have an offset that isn't handled.  Ooops. 


Trying Q's recipe for engaging AO path

I am able to get the MC2 AO gain slider up to -10dB (-7 is also okay).  If I increase the digital CARM gain too much, I see gain peaking at about 800Hz, so something good is happening.  (That was with a CARM_B gain of 2.0 and CARM_A gain of 0.  Don't go to 2.0)

I tried once without engaging his 300:80 1/f^2 filter in the CM_SLOW filter banks to start stepping up the CM REFL1 and MC AO gains together, but I only made it 2 steps of 1dB each before I lost lock. 

I tried once or twice turning on that 300:80 filter that Q said over the phone really helped his PRY locking, but it causes loop oscillations in CARM.  Also, I forgot to turn it off for ~45 minutes, and it caused several locklosses.  Ooops.  Anyhow, this isn't the right filter for this situation.


AS55 whitening problem

Twice I tried turning on the AS55 whitening.  Once, I was only partly transitioned from ALSdiff to AS55, the other time was the one time I made the full transition.  It caused the lockloss from the only RF-only lock I had tonight :(

Unfortunately I don't have the time series before the whitening filters (not _DQ-ed), but you can see a giant jump in the _ERR signals when I turn on the whitening, just before the arm power dies:

AS55whitening_lockloss_12March2015.pdf

The AS55 phase is -30, I has an offset of 28.2 and Q has an offset of 6.4.  Both have a gain of 1.  This should give us enough info to back out what the _IN1 signals looked like before I turned on the whitening if that's useful.


Other random notes

Ramp times for CARM_A, CARM_B, DARM_A and DARM_B are all 5 seconds.  This is set in the carm_cm_up script.

carm_cm_up script freezes the arm ASS before it starts the IR->ALS transition, to make it more convenient to run the ASS each lockloss.

carm_cm_up script no longer has a bunch of stuff at the bottom that we're not using.  It's all archived in the svn, but the remnants from things like variable finesse aren't actively  useful.

carm_cm_down script turns off the CM_SLOW whitening (which gets set in the up script)

carm_cm_down script clears the history of the ETM oplevs, in case they went bad (from some near divide-by-zero action?), but the watchdog isn't tripped. This clears away all the high freq crap and lets them do their job.

FSS Slow has been larger than 0.55 all night, larger than 0.6 most of the night, and larger than 0.7 for the last bit of the night.  MC seems happy.

both carm_cm_up and carm_cm_down are checked into the svn.  The up script is rev 45336 and the down script is 45337.

Some offset (maybe the fact that the fast AO path had an un-compensated offset?) is pulling the arm powers down as I make the transitions:


Recipe overview

  • Lock PRMI with arms held on ALS at 3nm CARM offset.  Bring CARM offset to 0.
  • Turn on CARM_B and DARM_B a little bit, then turn on their integrators
  • Lower the PRCL and MICH gains a little.
  • Increase the CARM_B gain a bit, then turn off FM1 for both CARM and DARM.
  • Increase CARM_B gain, lowering CARM_A gain.
  • Increase DARM_B gain, lowering DARM_A gain.  Now the power should definitely be stable (usually ends up around 80).
  • Partly engage AO path.
    • CM board REFL1 gain = -20dB
    • CM board AO gain = 0dB
    • MC2 board AO gain starts at -32dB, stepped up to -20dB
  • Increase CARM_B gain a bit
  • More AO path:  MC2 board AO gain steps from -20dB to -10dB
  • Increase CARM_B gain to 1.5, turn CARM_A gain to zero
  • CM_SLOW whitening on

After that, I by-hand made the DARM transition on the 6th successful scripted CARM transition, and tried to script what I did, although I was never able to complete the DARM transition again.  So, starting where the recipe left off above,

  • Turn off DARM's FM2 boost to win some more phase margin.
  • Increase DARM_B gain to 0.5, lower DARM_A gain to 0.

Since DARM doesn't have an analog fast path, it is stuck in the delicate filter situation.  I think that I should probably start using the UGF servo once the arm power is stable so that DARM stays in the middle of its phase bubble.

Rather than typing out the details of the recipe, I am attaching the up script.

Attachment 1: AS55whitening_lockloss_12March2015.pdf  16 kB  | Hide | Hide all
AS55whitening_lockloss_12March2015.pdf
Attachment 2: MoreDARMB_powerWentDown_12March2015.png  504 kB  | Hide | Hide all
MoreDARMB_powerWentDown_12March2015.png
Attachment 3: carm_cm_up_zip.sh.gz  1 kB
ELOG V3.1.3-